the density of N2 gas at 10C and 1.5 atm is
a) 2.8g/mL b)18g/L c)51.2g/l d) 1.8g/l e)28g/l
Answer #1

Answer -

Temperature = 10\degree C or 283.15 K [°C + 273.15 = K]

Pressure = 1.5 atm

Molar Mass of N2 = 28 g/mol

Density of N2 gas = ?

We know that,

PV = nRT

where,

n = mass / Molar Mass

R = Gas Constant

T = Temperature

P = Pressure

PV = Mass*R*T/Molar Mass

(P*Molar Mass) / (R*T) = (Mass/Volume) -----1

Also,

Density = Mass /Volume -----2

From 1 and 2,

So,

Density = (P*Molar Mass) / (R*T)

Put the values,

Density = (1.5 atm * 28 g/mol) / (0.08206 L⋅atm⋅K−1⋅mol−1 * 283.15 K)

Density = 1.81 g/mL [ANSWER]